# Empirical Analysis of Pull Requests for Google Summer of Code

### v1

@@ Abstract
Internship and industry-affiliated capstone projects are popular ways to expose students to real world experience and bridge the gap between academic training and industry requirements. However, these two approaches often require active industry collaboration and many students often struggle to find industry placements. Open-source contributions is a crucial alternative to gain real world experience, earn publicly verifiable contribution with real world impact, and learn from experienced open-source contributors. The Google Summer of Code is a global initiative that matches students or new contributors with experienced mentors to work on open-source projects. The goal of the program is to introduce the students to open-source, help gain valuable skills under the guidance of a mentor, and hopefully continue to contribute to open source development; thereby, provide a continuous pool of talented new contributors necessary for maintaining an open source project. This study presents an empirical analysis of pull requests created by interns during the Google Summer of Code program. We extracted and analysed 17,232 pull requests from 2456 interns across 1937 open-source projects. The results show that majority of the tasks involve both code-intensive tasks like adding new features and fixing bugs as well as non-code tasks like updating the documentation and restructuring the code base. The feedback from reviewers covers code functionality and programming logic, testing coverage, error handling, code readability, and adopting best practices. Finally, we discuss the implications of these results for software engineering education.
@@ I Introduction
Majority of the software engineering education research are usually directed towards students within formal academic setting. Unfortunately, these traditional approaches to software engineering education often limit students’ exposure and engagement to real-world projects; thereby, failing to fully harness their potential and creativity [1],[2]. The industry has often noted that fresh graduates have limited exposure to real-world projects and are often ill-equipped to deal with industrial projects [3],[4],[5],[6]. A number of approaches such as internships and university-industry collaboration on capstone projects have been designed to expose students to real-world projects and provide valuable industrial experience. However, these initiatives often require an active industry partner, and this may not be feasible in all places at all times.
@@
In the same vein, open-source software communities are continuously in need of new contributors to maintain and sustain the code base. Contributions to open source can be publicly verified and may be used to demonstrate competence during a job search. Furthermore, open-source code repositories are abundant, easy to access, and provide real-world examples that are needed to prepare students for the industry. Therefore, recruiting students to work on open-source code repositories will equip students with the necessary industrial experience, provide publicly-verifiable evidence of competence that can be useful for students during a job search, and help with the needed pool of contributors necessary to maintain an open-source code base.
@@
The Google Summer of Code (GSoC) [^1] https://summerofcode.withgoogle.com/[7] is an annual, global three-month internship program that connects students to open-source software communities. The program provides students the opportunity to work on real-world open-source projects, with guidance from experienced mentors in the field. Google provides stipends to student participants after completing certain milestones. The program is the first global initiative to introduce students to open source software development and bring in new contributors. Since 2005, the program has connected more than 19,000 new contributors from 112 countries with 18,000 mentors from 133 countries. The program has generated over 43 million lines of code across numerous projects for more than 800 open-source organizations [^2] https://summerofcode.withgoogle.com/about. The code contributions of the participants are publicly available. Therefore, the GSoC program can provide valuable data to gather insights into activities of novices in open-source development.
@@
Contributions to open-source are often done via pull requests where a contributor first clone the code base, add changes (e.g, add a new feature) to the local copy, and then notifies the repository team that they have completed a feature or modified the code base [8]. Collaborators can then review and provide comments on the proposed changes. Code review comments provide valuable feedback that guide improvements in the software development process. A maintainer has to approve the requested change before they can be integrated into the main code base. Although a number of studies have been conducted on pull requests, there is limited research on the analysis of pull requests by interns or students such as pull requests of participants in the Google Summer of Code program.
@@
This study provide an analysis of pull requests created by interns during the Google Summer of Code program from 2020 to 2023 (a 4-year period). We analyze pull request data, review comments, project-related features and contributor (intern) data to provide informative insights into the kind of contributions the intern’s make, the nature of feedback they receive, and possible implications for software engineering education. The study analyzed 17232 pull requests by 2456 interns across 1937 projects. We then used topic modelling techniques to gain insight into the common themes in the pull request title, description and review comments. In summary, the paper provides the following contributions
@@
    A publicly available dataset  [^3] https://github.com/compedutech/gsoc of 17,232 pull requests created by GSOC interns from 2000 to 2023. These pull requests were extracted directly from Github via the GiHub API. The scripts used for web scrapping (to identify the GitHub URL) and the script for extracting the pull requests from GitHub are also provided.
@@
    An insight into the success or failure of the GSOC program in promoting open-source contribution. Specifically, we provide information on further contributions by the interns to the same project after the GSOC program.
@@
    An analysis into the type of task the interns worked on and the feedback they receive from the other project collaborators.
@@
    A discussion on the possible implication of the study results for software engineering education.
@@ II Background on Google Summer of Code Program
Google Summer of Code (GSoC) is an annual, global program run by Google that encourages university students or new contributors to participate in open-source software development. The program pairs new contributors with mentors from open-source organizations. Google provide stipends to contributors who successfully meet some specified milestones. The following paragraph provides more detailed overview of the program.
  • Goal. GSoC is designed to introduce students to open-source software development by working on real-world projects with guidance from experienced mentors. The program has two main goals. Firstly, to expose students or novice to open-source development under the guidance of experience mentors and gain real-world experience. Secondly, to provide open-source organizations with new contributors that can continue to contribute to the open source communities long after the program is completed.
  • Eligibility. Participants must be students or new to open-source from all over the world. They must be atleast 18 years old, eligible to work in their country of residence, and residence country should not be currently embargoed by the United States.
  • Organization. Open-source organizations apply to be mentoring organizations, offering projects for students to work on. The organization must have an active open source project and not be based in a currently embargoed by the United States.
@@
  • Timeframe. The standard timeline for the program is 12 weeks and it usually runs over the summer ( from May to August). However, the mentors and contributors can agree to extend the program till 22 weeks.
  • Application process. First, open-source organizations submit proposals for specific software development projects. Students would then create proposals explaining how they would approach the project. The mentoring organizations would then select students based on their proposals and the skills demonstrated.
  • Onboarding and mentorship. Selected students are paired with mentors from the open-source organization. The student spend the first few weeks to get familiar with the code base and work on the expected milestones in collaboration with their mentors. The students then spend the next twelve (or more) weeks working on the implementation of the proposed solution and completion of the milestone.
  • Completion. Students work on their projects throughout the summer, delivering code, features, or bug fixes, and submitting reports to both their mentors and Google.
  • Stipend. Google provides stipends to students based on their location and the duration of the program. The payment of the stipend are often contingent on the completion of some specified milestones.
@@
    Career development. The contributions implemented during the GSoC program are public and easy to verify. Many students saw the GSoC program as an opportunity to build their resume, kickstart their career, and increase their visibility when seeking a job [9]. Furthermore, the program is popular, highly regarded by employers, and many students go on to work for the organizations they contributed to. This makes participation in the program an excellent component of the participants resume when searching for jobs.
@@
    Open-source sustainability. The program provides new contributors to open-source communities which is necessary for maintaining the open-source projects. Tan et al. [10] noted that sustainability of the project is a major motivating factor for mentors to participate in the program.
The GSoC program offers an excellent avenue to introduce students or newcomers to open source, and support the sustainability of open-source communities. Majority of the organizations participating in the program have also reported that the program helps them to find new contributors who are active and committed to the open source project [11]. The contributions of the students to the open source community are also public and accessible to anyone. This public data provides valuable dataset for analyzing the impact of the program, understand the challenges or barriers to open source software development, and probe factors that ensure successful engagement with the open source communities.
@@ III Literature Review
A number of research has been conducted on summer code programs and pull requests, but to the best of our knowledge, none has provided empirical analysis of the pull requests developed during these summer programs. Silva et al. [9][12] conducted a qualitative study to understand what motivated students to participate in the GSoC program and then developed a theoretical framework to model student engagement. Their findings show that students often participate in the program to gain new software engineering skills, enhance their resume, and earn stipends. Tan et al. [10] also conducted a qualitative study on mentor’s participation in the GSoC program. The results of their analysis shows that intristic motivation to sustain the open source code base, gain new skills, and increased recognition in the open-source community are the major reasons why mentors participate in the program. Trainer et al. [13] conducted a study of 22 GSoC projects and 22 hackatons to understand student engagements with the open-source scientific community. Their findings reveal that private communications with the mentor often create strong ties with the mentor while public sharing of progress report creates weak ties with other community members. They also discovered that about 18% of the students later became mentors for others.
@@
Majority of the approach so far have used qualitative study (without considering the code base) to analyze students and mentors participation in the GSoC program. The closest to our work was conducted by Silva et al. [14][15] who analyzed the code commits of 866 students in the GSoC program from 2013 to 2015. The results show that their is a strong correlation between the number of commits a student produce and the retention of that student to the project during and after the GSoC program. They also noted that 82% of the projects merged atleast one commit of the students to their code base. Commits provide low-level data on the students contribution as they provide information on how and why each line of code was changed. Hence, it may be challenging to analyse the kind of tasks or features they work on. Pull requests on the other hand, contains aggregate changes from multiple commits and provide a high-level data on the kind of tasks the students work on. Pull request also provide opportunities for other members of the community to review the changes, provide comments or feedback, and suggest modifications to the proposed changes [8]. Furthermore, pull requests have been shown to be more precise than commits in evaluating the contributions of community members [16]. Our study extend the work by Silva et al. [14][15] by using pull requests to understand student’s contribution to the GSoC program.
@@
There has also been studies conducted on pull requests. Gousios et al. [8] conducted an exploratory study of the pull based model on 291 projects. They found out that the pull based model reduces the time needed to merge contributions to the code base, increases the development turnover, and provided more opportunities for community members to engage each other. Kalliamvakou et al. [17] also conducted a study on GiHub repositories and found out that pull requests are often used for code reviews and issue resolution. liu et al [18] also noted that Pull request can lead to more active code repositories. Yu et al. [19] and Kerzazi et al. [20] developed theoretical frameworks for suggesting reviewers to review pull requests. Zhang et al. [21] examined competing pull requests and noted that 31% of pull requests in 75% of the repositories they studied tend to modify the same code during the same period. Jiang et al. [22] studied reopened pull requests and noted that reopened pull requests often had more comments and lower rate of being merged with the code base than pull requests that have not been reopened. Legay et al [23] examined the impact of the acceptance or rejection of a pull request on future contribution to the project. Their results show that contributors with higher acceptance rate are more likely to continue contributing to the project while those with rejections are less likely to contribute in the future. The studies discussed in this paragraph have examined pull requests in different context. However, none of the studies have focused on pull requests produced by novice programmers or students. This study extends the literature to provide an in-depth analysis of pull requests from students or new contributors in the GSoC program.
@@
Internships and industry-based capstone projects are a popular way of providing students with real world experience. There has been many studies conducted within the context of internships and capstone projects. Kapoor et al. [24] and Lehman et al. [25] reported that internships often lead to improved professional development, while Groeneveld et al [26] noted that both internships and capstone projects promotes the development of soft skills among students. Dean et al . [27] noted that the impact of internships on students depend on the role of the student during project execution. Sudol et al. [28] also reported that students who had completed at least one internships tend to have reduced misconceptions about software engineering concepts. The related work on internships discussed so far have adopted qualitative studies to analyze the impact of internships on students without looking at the actual code contributions. The closest work to our study was by Menezes et al. [29] where they described their experience on an internship model that pairs students with volunteer mentors in open source projects. The authors then conducted a qualitative study using surveys and reports to understand the impact of the internship model on the student skills and job placement. Their results show that significant increase in student’s technical skills, soft skills, and job placement after graduation. They also noted that students who are paid stipends are more willing to work full time on the project.
This literature review section have presented a number of works related to GSoC, pull request and internships. However, none of the work examined has conducted a large-scale study on pull requests within the context of internships or summer coding programs. This study extends the literature review by analyzing the pull requests of students or interns participating in the GSoC program.
@@ IV Design Methodology
This section discusses the data collection process, the research methodology adopted in this paper, and the research questions we aim to answer.
@@ IV-A Data Collection
We extracted pull request data via web scrapping and the GitHub API. The following paragraphs highlights how we extracted the needed data.
@@
    Firstly, we used the Selenium Python Package to scrape the URLs of each intern’s personal reflection website from the GSOC web page for each year (2020, 2021, 2022 and 2023).
@@
    Secondly, we extracted all the URL links from each intern’s webpage (by searching for ”a” html tag).
@@
    Thirdly, we filtered out the links containing ”github” and (”issues” or ”pull”). For example the url link ”https://github.com/repos/asyncapi/community/pulls/720” contains the words ”github” and ”pull”. Most pull requests url on github often follow this pattern and we believe is the easiest way to extract most of the pull request link.
@@
    Finally, we used the GitHub API to extract the pull request data associated with the filtered links.
We extracted three main types of data related to the pull request, the contributor (intern) that created the pull request, and the project (or code repository) associated with the pull request. At the end of the Data collection process, we collected 17,232 pull requests from 2080 interns across 1937 open-source projects. The following subsections provides statistical discussion of the pull requests, projects and users embedded in the data.

### v2

@@ Abstract
Internship and industry-affiliated capstone projects are popular ways to expose students to real world experiences and bridge the gap between academic training and industry requirements. However, these two approaches often require active industry collaboration, and many students struggle to find industry placements. Open-source contributions are a crucial alternative to gain real world experience, earn publicly verifiable contribution with real-world impact, and learn from experienced open-source contributors. The Google Summer of Code (GSoC) is a global initiative that matches students or new contributors with experienced mentors to work on open-source projects. The program aims to introduce the students to open-source development, help them gain valuable skills under the guidance of mentors, and hopefully encourage them to continue contributing to open-source projects. The realization of the program objectives will provide a continuous pool of talented new contributors necessary for maintaining open-source projects. This study presents an empirical analysis of pull requests created by interns during the GSoC program. We extracted and analyzed 17,232 pull requests from 2,456 interns across 1,937 open-source projects. The results show most tasks involve both code-intensive activities like adding new features and fixing bugs, as well as non-code tasks like updating documentation and restructuring the codebase. Feedback from reviewers covers code functionality and programming logic, testing coverage, error handling, code readability, and adherence to best practices. Finally, we discuss the implications of these results for software engineering education.
@@ I Introduction
Most software engineering education research targets students in formal academic settings. Unfortunately, traditional approaches to software engineering education in these academic settings often limit students’ exposure to and engagement with real-world projects, failing to fully harness their potential and creativity [1],[2]. The industry frequently notes that fresh graduates have limited exposure to real-world projects and are often ill-equipped handle industrial challenges [3],[4],[5],[6]. Initiatives such as internships and university-industry collaborations on capstone projects have been designed to provide students with valuable industrial experience. However, these initiatives often require active industrial partnerships, which may not be always feasible.
@@
Open-source software communities continuously need new contributors to maintain and sustain their codebase. Contributions to open source can be publicly verified and may demonstrate competence during job searches. Open-source code repositories are abundant, easily accessible, and provide real-world examples that are essential to prepare students for the industry. Therefore, recruiting students to work on open-source code repositories equips the students with industrial experience, publicly-verifiable evidence of competence, and helps sustain the pool of contributors needed for open-source projects.
@@
The Google Summer of Code (GSoC) [^1] https://summerofcode.withgoogle.com/[7] is an annual, global, three-month internship program that connects students to open-source software communities. Participants work on real-world open-source projects under the guidance of experienced mentors. Google provides stipends to student participants after completing certain milestones. The program is the first global initiative to introduce students to open source software development and bring in new contributors. Since 2005, GSoC has connected more than 19,000 new contributors from 112 countries with 18,000 mentors from 133 countries. The program has generated over 43 million lines of code across numerous projects for more than 800 open-source organizations [^2] https://summerofcode.withgoogle.com/about. The code contributions of the participants are publicly available, thereby providing valuable data to study novice activities in open-source development.
@@
Contributions to open-source are often made via pull requests where contributors clone a codebase, make changes (e.g, add a new feature) locally, and then notify the repository team that they have completed a feature or modified the codebase [8]. Collaborators then review and provide comments on the proposed changes. Code review comments provide valuable feedback that guide the contributors and enhance code quality. Maintainers must approve the requested changes before they can be integrated into the main codebase. Although several studies have examined pull requests, there is limited research on the analysis of pull requests created by interns or students such as GSoC participants.
@@
This study analyzes pull requests created by interns during the GSoC program from 2020 to 2023 (a 4-year period). We examine pull request data, review comments, project-related features, and contributors (intern) data to gain informative insights into the contributions interns make, the feedback they receive, and possible implications for software engineering education. The study analyzed 17,232 pull requests by 2,456 interns across 1,937 projects. We used topic modeling techniques to gain insights into common themes in the pull request titles, descriptions and review comments. In summary, the paper provides the following contributions:
@@
    A publicly available dataset [^3] https://github.com/compedutech/gsoc of 17,232 pull requests created by GSOC interns from 2000 to 2023. This dataset was extracted directly from the official GSoC program page [^4] https://summerofcode.withgoogle.com (via web scrapping) and Github (via the GiHub API). The dataset also include the scripts used for web scrapping and extracting the pull requests from GitHub.
@@
    Insights into the success or failure of the GSOC program in promoting open-source contributions. Specifically, we provide information on interns’ continued contributions to the same project after the program.
@@
    An analysis of the tasks interns worked on and the feedback they received from the project collaborators.
@@
    A discussion on the implications of these results for software engineering education.
@@ II Background on Google Summer of Code Program
Google Summer of Code (GSoC) is an annual, global program run by Google that encourages university students or new contributors to participate in open-source software development. The program pairs new contributors with mentors from open-source organizations. Google provide stipends to contributors who successfully meet some specified milestones. The following paragraphs provide a detailed overview of the program.
  • Goal. GSoC aims to introduce students to open-source software development by engaging them in real-world projects under the guidance of experienced mentors. The program has two main goals. Firstly, to expose students or novices to open-source development and help them gain real-world experience. Secondly, to provide open-source organizations with new contributors that will continue to contribute to open-source communities even after the program is completed.
  • Eligibility. Participants must be students or new to open-source, at least 18 years old, eligible to work in their country of residence, and not residing in countries currently embargoed by the United States. The program is open to participants from all over the world who meet the criteria stated above.
  • Organization. Open-source organizations apply to mentor students, and offer projects for students to work on. Organizations must have active open-source projects and not be based in embargoed countries.
@@
  • Timeframe. The standard timeline for the program is 12 weeks and it usually runs from May to August. However, mentors and contributors may agree to extend the program to 22 weeks.
  • Application process. First, open-source organizations submit proposals for specific software development projects. Students would then create proposals explaining how they would approach the project. The mentoring organizations would then select students based on their proposals and demonstrated skills.
  • Onboarding and mentorship. Selected students are paired with mentors from the open-source organization. The student spend the first few weeks to get familiar with the codebase and collaborate with mentors on expected project milestones. The students then spend the next twelve (or more) weeks working on implementing solutions and completing milestones.
  • Completion. Students deliver code, features, or bug fixes, and submit reports to both their mentors and Google.
  • Stipend. Google provides stipends to students based on their location and duration of the program. The payment of the stipend are often contingent upon completion of some specified milestones.
@@
    Career development. The contributions implemented during the GSoC program are public and easy to verify. Many students saw the GSoC program as an opportunity to build their resume, kickstart their career, and increase their visibility when seeking a job [9]. Furthermore, the program is popular, highly regarded by employers, and many students go on to work for the organizations they contributed to. This makes participation in the program an excellent component of participants’ resume when searching for jobs.
@@
    Open-source sustainability. The program provides new contributors to open-source communities which is necessary for sustaining the open-source projects. Tan et al. [10] noted that sustainability of the project was a major motivating factor for mentors to participate in the program.
The GSoC program offers an excellent avenue to introduce students or newcomers to open source, and support the sustainability of open-source communities. Majority of the organizations participating in the program have also reported that the program helps them to find new contributors who are active and committed to the open source project [11]. The contributions of students to the open source community are also public and accessible to anyone. This public data from students’ contributions provides a valuable dataset for analyzing the program’s impact, understanding challenges or barriers to open-source software development, and probe factors that ensure successful engagement with the open source communities.
@@ III Literature Review
Several studies have examined the impact of summer coding programs, internships, and pull requests on software development and education. However, to the best of our knowledge, none of the studies have provided empirical analysis of pull requests created during summer programs. Silva et al. [9][12] conducted a qualitative study to understand students’ motivations for participating in GSoC program, and they then developed a theoretical framework to model students’ engagement. Their findings reveal that students often joined the program to acquire new software engineering skills, enhance their resumes, and earn stipends. Tan et al. [10] explored mentors’ motivations in the GSoC program and noted that intrinsic motivation to sustain the open-source codebase, learn new skills, and increased recognition in the open-source community are the major reasons why mentors participate in the program. Trainer et al. [13] analyzed 22 GSoC projects and 22 hackatons to understand student engagements with the open-source scientific community. Their findings reveal that private mentor-student communications often create strong ties, while public sharing of progress reports create weak ties with other community members. Interestingly, 18% of students later became mentors.
@@
Majority of the approach so far have used qualitative study (without considering the codebase) to analyze students and mentors participation in the GSoC program. The closest to our work was conducted by Silva et al. [14][15] who analyzed the code commits of 866 students in the GSoC program from 2013 to 2015. The authors discovered a strong correlation between the number of commits and students’ retention within projects during and after the GSoC program. Additionally, 82% of the projects merged at least one commit from students into their codebases. However, while commits provide granular insights into code changes, they may not fully capture task types or objectives. Pull requests on the other hand, contains aggregate changes from multiple commits and provide a high-level data on the kind of tasks the students work on. Pull requests also provide opportunities for other members of the community to review changes, provide comments or feedback, and suggest modifications to the proposed changes [8]. Furthermore, pull requests have been shown to be more precise than commits in evaluating the contributions of community members [16]. Our study extends the work by Silva et al. [14][15] to use pull requests for understanding students’ contributions to the GSoC program.
@@
Studies on pull requests have revealed their significance in collaborative software development. Gousios et al. [8] conducted an exploratory study of the pull-based model on 291 projects. They found out that the pull-based model reduces the time needed to merge contributions to the codebase, enhances development turnover, and encouraged community engagement. Kalliamvakou et al. [17] also conducted a study on GitHub repositories and noted that pull requests are often used for code reviews and issue resolutions. Liu et al [18] reported that pull requests can lead to more active code repositories. Yu et al. [19] and Kerzazi et al. [20] developed theoretical frameworks for recommending pull requests’ reviewers. Zhang et al. [21] examined competing pull requests and noted that 31% of pull requests in 75% of repositories tend to modify the same code during the same period. Jiang et al. [22] examined reopened pull requests and noted that reopened pull requests often had more comments and lower rate of being merged with the codebase. Legay et al. [23] examined the impact of the acceptance or rejection of a pull request on future contributions to a project. Their results show that contributors with higher acceptance rate are more likely to continue contributing to the project, while those with rejections are less likely to contribute in the future. The studies discussed in this paragraph have examined pull requests in different contexts. However, none of the studies have examined pull requests created by novice programmers or students. This study extends the literature to provide an in-depth analysis of pull requests from students or new contributors in the GSoC program.
@@
Internships and industry-based capstone projects play crucial roles in providing students with real-world exposure. Kapoor et al. [24] and Lehman et al. [25] found that internships significantly improves professional development, while Groeneveld et al. [26] noted that both internships and capstone projects promotes the development of soft skills among students. Dean et al. [27] emphasized that the impact of internships on students varies by student involvement levels. Sudol et al. [28] reported that internship experience reduces misconceptions about software engineering concepts. The related work on internships discussed so far have adopted qualitative studies to analyze the impact of internships on students without considering the actual code contributions. The closest work to our study was by Menezes et al. [29] where they described their experience on an internship model that pairs students with volunteer mentors in open-source projects. The authors then conducted a qualitative study using surveys and reports to understand the impact of the internship model on students’ skills and job placement. Their results show significant increase in students’ technical skills, soft skills, and job placement after graduation. They also noted that students who are paid stipends are more willing to work full time on the project.
This literature review section have discussed studies related to GSoC, pull requests and internships. Although, these studies provided valuable insights, none of them conducted large-scale analyses of pull requests within the context of internships or summer coding programs. This study fills this gap by examining GSoC participants’ pull requests, to offer a comprehensive view of their contributions and feedback.
@@ IV Design Methodology
This section outlines the data collection process, research methodology, and the research questions we aim to answer.
@@ IV-A Data Collection
We extracted pull request data using web scrapping and the GitHub API. The following paragraphs highlights how we extracted the needed data.
@@
    The Selenium Python Package was used to scrape the URLs of each intern’s personal reflection website on the GSoC web page for the years 2020 to 2023.
@@
    The personal reflection websites of the GSoC participants often contained links to their contributions during the program. So, we extracted all URLs from each intern’s webpage by locating HTML anchor tags.
@@
    URLs containing ”GitHub” and either ”issues” or ”pull” were filtered to identify pull request links. For example the url link ”https://github.com/repos/asyncapi/community/pulls/720” contains the words ”github” and ”pull”. Most pull requests URLs on GitHub often follow this pattern and we believe it is the easiest way to extract most of the pull request links.
@@
    The GitHub API was used to extract the pull request data associated with the filtered URLs.
The dataset includes three main categories: pull request details, contributor (intern) information, and the project metadata. The final dataset comprises of 17,232 pull requests submitted by 2,456 interns across 1,937 open-source projects. The following subsections provides statistical overview of the pull requests, projects and users (contributor) information embedded in the dataset.